OpenStudy (anonymous):

4-x = l2-3xl pretend the l's are absolute value signs

OpenStudy (mathstudent55):

Let's say you have an absolute value equation of this type |X| = Y where X and Y are expressions. To solve it, get rid of the absolute value signs by solving this complex equation: X = Y or X = -Y

OpenStudy (mathstudent55):

@Meep. Do you understand the explanation above, or do you need more help?

OpenStudy (anonymous):

I think I need more help.

OpenStudy (mathstudent55):

Ok. That's why I'm here.

OpenStudy (anonymous):

Is there a specific step to getting rid of the sings or do you just pretend like they aren't there and solve it like a normal equation?

OpenStudy (anonymous):

*signs

OpenStudy (mathstudent55):

Notice the explanation. You have an expression inside absolute value signs equaling another expression. In your case, the expression 2 - 3x is inside the absolute value signs. The expression 4 - x is the second expression not in the absolute value signs.

OpenStudy (mathstudent55):

To get rid of the absolute value signs, you must solve two equations separated by the word "or."

OpenStudy (mathstudent55):

The first equation is simply the equation you have just dropping the absolute value signs.

OpenStudy (mathstudent55):

The second equation is similar, but when you drop the absolute value signs, you must put a negative sign and parentheses around the expression that was in absolute values signs.

OpenStudy (mathstudent55):

First, let me just do one step to make this easier to follow. Let's take your equation as it appears.

OpenStudy (mathstudent55):

\(4 - x = |2 - 3x|\)

OpenStudy (mathstudent55):

Let's do one step before we start solving. Let's switch sides. \(|2 - 3x| = 4 - x\)

OpenStudy (mathstudent55):

All I did was switch sides to have the absolute value part on the left side.

OpenStudy (mathstudent55):

Now we do the step of getting rid of the absolute value.

OpenStudy (mathstudent55):

\(2 - 3x = 4 - x\) or \(-(2 - 3x) = 4 - x\)

OpenStudy (anonymous):

so the answer should be x= -1 or x= 1 ?

OpenStudy (mathstudent55):

Now you need to solve both equations.

OpenStudy (mathstudent55):

2 - 3x = 4 - x or -(2 - 3x) = 4 - x -2 = 2x or -2 + 3x = 4 - x x = -1 or 4x = 6 x = -1 or x = 6/4 = 3/2 Answer is: x = -1 or x = 3/2

OpenStudy (anonymous):

Okay! I got it now. :) thanks for the help.

OpenStudy (mathstudent55):

You're welcome. The important thing here is for you to understand how to solve an absolute value equation.

OpenStudy (mathstudent55):

First, have the absolute value part on one side of the equal sign, and the rest on the other side. Then solve the following two equations separated by the word "or." The first equation is simply the equation with the absolute value signs dropped. The second equation is: drop the absolute value signs and place what was inside the absolute value signs inside parentheses and a negative sign outside.